# Efficiently Identifying Watermarked Segments in Mixed-Source Texts

### v1

@@ A.1 FPR Calibration Function $F$
As discussed in Section 3.2, the FPR Calibration Function calculates the p-value / FPR for per-instance watermark detection, given the detection scores and the original text. We follow the methodologies outlined in Zhao et al. [2023] and Fernandez et al. [2023] for FPR calibration. This section presents three methods for detecting KGW-Watermark, Unigram-Watermark, and Gumbel-Watermark, each employing a unique scoring mechanism and statistical test to assess the FPR.
A.1.1 KGW-Watermark
@@
For the KGW-Watermark scheme described in Kirchenbauer et al. [2023], we follow the approach in Fernandez et al. [2023]. When detecting the watermark for a text segment, under the null hypothesis $\mathcal{H}_{0}$ (i.e., the text is not watermarked), the score $\operatorname{Score}(y)=\sum_{t=1}^{n}\mathbf{1}(y_{t}\in\text{Green Tokens})$ follows a binomial distribution $\mathcal{B}(n,\gamma)$, where $n$ is the total number of tokens and $\gamma$ is the probability of a token being part of the green list. The p-value for an observed score $s$ is calculated as:
@@
A.1.2 Unigram-Watermark
For the Unigram-Watermark scheme, we adopt the methodologies from Zhao et al. [2023]. To achieve a better FPR rate, the detection score differs from the KGW-Watermark approach. The score is defined as $\operatorname{Score}(y)=\sum_{t=1}^{m}\mathbf{1}(\tilde{y}_{t}\in\text{Green Tokens})$, where $\tilde{y}=\text{Unique}(y)$ represents the sequence of unique tokens in text $y$, and $m$ is the number of unique tokens.
@@ A.1.2 Unigram-Watermark
$\mathsf{Var}\left[\sum_{t=1}^{m}\mathbf{1}(\tilde{y}_{t}\in\text{Green Tokens})\mid y\right]=m\gamma(1-\gamma)(1-\frac{m-1}{n-1}),$
@@
A.1.3 Gumbel Watermark
For the Gumbel Watermark [Aaronson, 2023], we adopt the approach presented in Fernandez et al. [2023], which utilizes a gamma test for watermark detection. Under the null hypothesis $\mathcal{H}_{0}$, $\operatorname{Score}(y)=\sum_{t=1}^{n}\log\left(1/(1-r_{y_{t}})\right)$ follows a gamma distribution $\Gamma(n,1)$. The p-value for an observed score $s$ is calculated as:
@@ A.1.3 Gumbel Watermark
A.2 Discussion and Limitation
Our methods, GCD and AOL, can be applied to other watermarking schemes as long as they have token-wise detection scores for the sequence, such as Hu et al. [2024] and Zhao et al. [2024]. The detection results are constrained by the strength of the original watermark generation and the quality of the prompt text. In some cases, low-quality text produced by the watermark generation method cannot be directly detected using the original detection method. Additionally, positive samples created by inserting the generated watermark paragraph into natural text may not be detectable with our approach. However, these limitations arise from the current limitations of watermark generation and detection methods themselves, which is outside the scope of detecting small watermarked segments within long text, the focus of this work. Therefore, we assume that our method needs only to detect reasonably high quality watermarked text segments embedded in long text.

### v2

@@ Appendix B FPR Calibration Function $F$
As discussed in Section 3.1, the FPR Calibration Function calculates the p-value / FPR for per-instance watermark detection, given the detection scores and the original text. We follow the methodologies outlined in Zhao et al. (2023) and Fernandez et al. (2023) for FPR calibration. This section presents three methods for detecting KGW-Watermark, Unigram-Watermark, and Gumbel-Watermark, each employing a unique scoring mechanism and statistical test to assess the false positive rate.
B.1 KGW-Watermark
@@
For the KGW-Watermark scheme described in Kirchenbauer et al. (2023), we follow the approach in Fernandez et al. (2023). When detecting the watermark for a text segment, under the null hypothesis $\mathcal{H}_{0}$ (i.e., the text is not watermarked), the score $\operatorname{Score}(y)=\sum_{t=1}^{n}\mathbf{1}(y_{t}\in\text{Green Tokens})$ follows a binomial distribution $\mathcal{B}(n,\gamma)$, where $n$ is the total number of tokens and $\gamma$ is the probability of a token being part of the green list. The p-value for an observed score $s$ is calculated as:
@@
B.2 Unigram-Watermark
For the Unigram-Watermark scheme, we adopt the methodologies from Zhao et al. (2023). To achieve a better FPR rate, the detection score differs from the KGW-Watermark approach. The score is defined as $\operatorname{Score}(y)=\sum_{t=1}^{m}\mathbf{1}(\tilde{y}_{t}\in\text{Green Tokens})$, where $\tilde{y}=\text{Unique}(y)$ represents the sequence of unique tokens in text $y$, and $m$ is the number of unique tokens.
@@ B.2 Unigram-Watermark
$\mathsf{Var}$ $\left[\sum_{t=1}^{m}\mathbf{1}(\tilde{y}_{t}\in\text{Green Tokens})\mid y\right]$
$=m\gamma(1-\gamma)(1-\frac{m-1}{n-1}),$
@@
B.3 Gumbel Watermark
For the Gumbel Watermark (Aaronson, 2023), we adopt the approach presented in Fernandez et al. (2023), which utilizes a gamma test for watermark detection. Under the null hypothesis $\mathcal{H}_{0}$, $\operatorname{Score}(y)=\sum_{t=1}^{n}\log\left(1/(1-r_{y_{t}})\right)$ follows a gamma distribution $\Gamma(n,1)$. The p-value for an observed score $s$ is calculated as: